OpenStudy (anonymous):

Find the angle between the vectors u and v if u = (- 1, 1) and v = (3, 1). Round answer to two decimal places. A. 114.64o B. 115.33o C. 116.57o D. 117.53o E. 117.98o

OpenStudy (amistre64):

cos(t) = v.u/|v||u| t = cos^-1(v.u/|v||u|)

OpenStudy (anonymous):

The angle between a horizontal line. ( say the x axis) and 1,1 is 45 The angle between the horizontal and 3,1 is 15.

OpenStudy (anonymous):

So the angle should be 30

OpenStudy (amistre64):

the angle is measured in counter clockwise rotation

OpenStudy (anonymous):

ok so amistre.... I have that formula and plugged in my numbers -2/llull llvll

OpenStudy (amistre64):

-1,1 3,1 ----- -3+1 = -2 = v.u

OpenStudy (anonymous):

right i got that

OpenStudy (amistre64):

|v| = sqrt(2); |u| = sqrt(10) = sqrt(20) = 2sqrt(5)

OpenStudy (amistre64):

-2 ------- = -1/sqrt(5) right? 2 sqrt(5)

OpenStudy (anonymous):

yes when simplified noew I take the arc cos of that

OpenStudy (anonymous):

*now

OpenStudy (anonymous):

so it's looking like c

OpenStudy (amistre64):

cos^-1(-1/sqrt(5)) = .... 116.5650

OpenStudy (amistre64):

C is it :)